« Project Status: 8 January 2008 2030UTC | Main | Project Status: 9 January 2008 2030UTC »

Project Status: 9 January 2008 0100UTC (IOP#5 Update)

Report from BLU

Contacted radar crew. Had a number of heavy snow bands, last one was 8/23-9/00UTC. Sergey estimated 9 new inches with a 6 to 1 snow to water ratio. Thus 1.5 inches of liquid.....looks like good event. Radar dish is now covered with 3-4mm of ice due to an encounter with supercooled cloud. DAX radar indicates showers moving eastward with back edge over Central Valley so they should end in a couple of hours. Heaviest are now east and south; only scattered showers remain near BLU. Surface front looks to be in Central Valley so a 9/02-03UTC passage looks good. Snow plow operator will plow road from 19-21L so radar crew can leave the site tonight. Plan is to shut down radar at 9/04UTC (20L) lock things up and head to Auburn.

McGinley, GSD

TrackBack

TrackBack URL for this entry:
http://www.etl.noaa.gov/cgi-bin/mt/mt-tb.cgi/714